OpenStudy (zubhanwc3):

a) f(x) = 4x - 9. Find K so that (integral of (f(t))dt + k) [-1, x] = the integral of f(t) dt from [3,x] b) find d/dx the integral f(t) dt from [-1,x]

ganeshie8 (ganeshie8):

looks pretty straightforward... have u tried anything yet ? :)

OpenStudy (zubhanwc3):

i have, but i dont know exactly what to do i tried plugging in x/-x and than taking the derivative, but i didnt know how that would help me find k

ganeshie8 (ganeshie8):

you have this for part \[\int\limits_{-1}^x f(t) dt + k = \int \limits_3^xf(t)dt\] ?

OpenStudy (zubhanwc3):

ya

ganeshie8 (ganeshie8):

familiar with below property ? \[\int\limits_a^c f(t) dt = \int\limits_a^b f(t) dt + \int\limits_b^cf(t) dt\]

OpenStudy (zubhanwc3):

ya

ganeshie8 (ganeshie8):

Using that property you can split the given equation \[\int\limits_{-1}^x f(t) dt + k = \int \limits_3^xf(t)dt \] like below \[\int\limits_{-1}^3 f(t) dt + \int\limits_{3}^x f(t) dt + k = \int \limits_3^xf(t)dt\] yes ?

OpenStudy (zubhanwc3):

wait, why isnt the bottom of the right side of the equation -1?

ganeshie8 (ganeshie8):

why do u think it has to be -1 ?

ganeshie8 (ganeshie8):

we are splitting ` -1 to x` into `-1 to 3` and `3 to x`

OpenStudy (zubhanwc3):

would the 3 on top and bottom cancel out?

ganeshie8 (ganeshie8):

nope, but u can cancel something really big.. look closer

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^3 f(t) dt + \color{red}{\int\limits_{3}^x f(t) dt} + k = \color{red}{\int \limits_3^xf(t)dt}\]

OpenStudy (zubhanwc3):

im confused

OpenStudy (zubhanwc3):

shouldnt the intervals be [-1,x] and what do you mean by something big is canceled out

ganeshie8 (ganeshie8):

[-1, x] changed to [-1, 3] + [3, x] using the property i gave you earlier

OpenStudy (zubhanwc3):

ya i understand that, but what you wrote with the intergral signs were different

ganeshie8 (ganeshie8):

im not sure what difference are u seeing :/

OpenStudy (zubhanwc3):

u put the intervals of [-1,3] + [3,x] = [3,x]

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

you're given the equation \[\int\limits_{-1}^x f(t) dt + k = \int \limits_3^xf(t)dt\]

ganeshie8 (ganeshie8):

splitting the integral on left side as [-1, 3] + [3, x] you get : \[\int\limits_{-1}^3 f(t) dt + \int\limits_{3}^x f(t) dt + k = \int \limits_3^xf(t)dt\]

OpenStudy (zubhanwc3):

oooo i get that now

ganeshie8 (ganeshie8):

okay good :)

OpenStudy (zubhanwc3):

it actually becomes [-1,x] but that = [3,x] from the original equation is that right?

ganeshie8 (ganeshie8):

yes

OpenStudy (zubhanwc3):

how do i use this to find k?

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^3 f(t) dt + \color{Red}{\int\limits_{3}^x f(t) dt} + k = \color{Red}{\int \limits_3^xf(t)dt}\] cancel that

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^3 f(t) dt + k = 0\] plugin the value of f(t) and evaluat ethe definite integral

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^3 (4t-9)~ dt + k = 0\]

ganeshie8 (ganeshie8):

\[k = -\int\limits_{-1}^3 (4t-9)~ dt \]

ganeshie8 (ganeshie8):

evaluate

OpenStudy (zubhanwc3):

aite, now how do i do b?

ganeshie8 (ganeshie8):

what do u get for k ?

OpenStudy (zubhanwc3):

-16

OpenStudy (zubhanwc3):

my calc says 20, did i mess up?

OpenStudy (zubhanwc3):

nvm, forgot to take the anti-derivative

ganeshie8 (ganeshie8):

20 is right

OpenStudy (zubhanwc3):

forgot to take anti-derivative , so i just plugged into 4x-9

OpenStudy (zubhanwc3):

yep i now got 20

ganeshie8 (ganeshie8):

sounds great!

ganeshie8 (ganeshie8):

for part b you want to compute \[\large \dfrac{d}{dx}\int\limits_{-1}^x f(t) dt \]

ganeshie8 (ganeshie8):

Say \(\int f(x) dx = F(x) + C \implies f(x) = F'(x)\) by fundamental thm of calculus \[\large \begin{align} \dfrac{d}{dx}\int\limits_{-1}^x f(t) dt &= \dfrac{d}{dx}\left[F(t) \right]_{-1}^x\\~\\ &=\dfrac{d}{dx}\left[F(x) - F(-1)\right]\\~\\ &=\dfrac{d}{dx}F(x) - \dfrac{d}{dx} F(-1)\\~\\ &= f(x) - 0\\~\\ &=4x-9 \end{align}\]

ganeshie8 (ganeshie8):

Notice that `F(-1)` is the value of function evaluated at x=-1 which will be a `constant`. so its derivative will be 0.

OpenStudy (zubhanwc3):

thank you, you helped me clear the only problem on the test review that i was stuck on